Solve the following initial value problem using the method of Laplace transform. y + 2y +10y = f(t); y(0)= 1, y(0) = 0, wh
Answer #1

Sofi- P Conceptit Let L {f(t)} = F(s) Then, Juse: Letit)} = sfcs) - f 10) L {f(t)} = 5² F (S) - sf (o)-fo) to ostdo 3 tilo ;Yeo) [s? 42s +1} - (5+2) = 10 cele 1] + 20 (o-e ] -105 -10 Vis [stas to] (s+2) = 10 - 10+ lue 5 S e s » Yes) [ 5² 425 +10] =(5+2) I sti x s coop - base] (st1²422 432 Stll 2422 txl sint T inst Result (oIl tem : SC 3725110) Apply Laplace inverse rule ft { fies)} = f ct) than I e a F(s)} = H (t-a) f (tra) where Hits in heavesicLes - To ten (54) 432 E l ) = certyf sin3t - Reant ② causine to results , . & D atlyco) h*(9) +CD) +2(T)+7%) y = {c* cost